OpenStudy (anonymous):

Find the exact length of the curve. Use a graph to determine the parameter interval. r = (cos(θ/2))^2

OpenStudy (anonymous):

aren't they suppose to tell you from where to where?

OpenStudy (anonymous):

You are suppose to find that out yourself by graphing it

OpenStudy (anonymous):

here is the graph but you cant use it tho to determine the interval

OpenStudy (anonymous):

https://www.google.com/#q=(cos(x%2F2))%5E2

OpenStudy (anonymous):

oh also is this in polar? if it is in polar then i am wrong lol

OpenStudy (anonymous):

ya it's polar

OpenStudy (anonymous):

use this polar grapher :)

OpenStudy (anonymous):

@timo86m Unfortunately, it's not a circle; it's a Cardioid. http://www.wolframalpha.com/input/?i=Polar+plot+%28cos%28t%2F2%29%29%5E2

OpenStudy (anonymous):

I would first rewrite the function as \(\large r=\cos^2(\theta/2)=\frac{1}{2}(1+\cos\theta).\) Then recall that arclength in polar coordinates is given by \[\large L = \int_{\theta_0}^{\theta_1} \sqrt{r^2+\left(\frac{dr}{d\theta}\right)^2}\,d\theta\] Although the wolfram alpha link shows the plot for \(\large 0\leq \theta\leq 4\pi\), it turns out that \(\large 0\leq \theta\leq2\pi\) is the smallest interval needed to plot the function. We now see that \[\large L= \int _0^{2\pi}\sqrt{r^2+\left(\frac{dr}{d\theta}\right)^2}\,d\theta\] where \(\large r=\frac{1}{2}(1+\cos(2\theta))\). Can you take things from here? :-)

OpenStudy (anonymous):

I get stuck at \[1/2\sqrt{(1+2\cos2\theta} +\cos(2\theta)^2 +4\sin (2\theta)^2)\]